40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7822     Entry time: Thu Dec 13 04:42:32 2012
Author: Den 
Type: Update 
Category: SUS 
Subject: ITMX local damping 

Tonight we've noticed that ITMX local damping was kicking the optics. This happened because LR shadow sensor was not working. In ~30 minutes it started to work again. Evan and I were working on installation, moving and focusing cameras and locking prcl and mich. We've installed a camera on BSC and plugged it in to PSL_SPARE input.

I'm not sure that this can be correlated to ITMX LR shadow sensor behaviour.

 

lrsen.png

ELOG V3.1.3-